# Fabien Despinoy

### v2

@@ Abstract
Semantic Segmentation is one of the most challenging vision tasks, usually requiring large amounts of training data with expensive pixel level annotations. With the success of foundation models and especially vision-language models, recent works attempt to achieve zeroshot semantic segmentation while requiring either large-scale training or additional image/pixel level annotations. In this work, we generate free annotations for any semantic segmentation dataset using existing foundation models. We use CLIP to detect objects and SAM to generate high quality object masks. Next, we build a lightweight module on top of a self-supervised vision encoder, DinoV2, to align the patch features with a pretrained text encoder for zeroshot semantic segmentation. Our approach can bring language-based semantics to any pretrained vision encoder with minimal training. Our module is lightweight, uses foundation models as the sole source of supervision and shows impressive generalization capability from little training data with no annotation.
@@ 1 Introduction
Large-scale and inexpensive training data has recently enabled the surge of foundation models in computer vision [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al.]. These models have been employed to avoid expensive annotations and computational requirements of many vision tasks [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al., Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al., Alayrac et al.(2022)Alayrac, Donahue, Luc, Miech, Barr, Hasson, Lenc, Mensch, Millican, Reynolds, et al.]. Leveraging foundation models for semantic segmentation requires the model to produce pixel-wise predictions generalizable to new datasets, domains and ontologies. Therefore, the model must be 1) promptable with an open set of categories and 2) highly discriminative for dense recognition tasks. Consequently, deploying foundation models to obtain cheap annotation for semantic segmentation is challenging as existing models lack either semantic awareness [Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Liu et al.(2022)Liu, Mao, Wu, Feichtenhofer, Darrell, and Xie, Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] or local feature robustness [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Jia et al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and Duerig].
[…]
In particular, there exists some evidence that Contrastive Language-Image Pretraining (CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.]), trained on a large set of image-text pairs, can be used to derive a semantic understanding of different regions within an image [Zhou et al.(2022)Zhou, Loy, and Dai, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.]. However, CLIP features are sub-optimal for dense recognition tasks such as semantic segmentation. On the other hand, self-supervised task-agnostic pretraining methods applied to a sufficiently large set of unlabelled image data have shown great success in producing robust and generic visual features, suitable for any vision task. Specifically, DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] features show generalization across different domains and are discriminative for dense recognition tasks. DINO models are however trained without any explicit notion of semantic meaning  [Zhang et al.(2022a)Zhang, Li, Liu, Zhang, Su, Zhu, Ni, and Shum, Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.].
@@
In this paper, we propose a novel approach to open vocabulary semantic segmentation by composition of different foundations models as building blocks and source of supervision. The result is a model that is highly discriminant, generalizable and grounded to semantic meaning and object shape without requiring human-generated segmentation or image annotations. We propose a lightweight alignment module which is trained in a contrastive manner to align visual features from a DINOv2 encoder with the text embedding space of a CLIP model. In order to supervise the shape and extension of objects, we generate pseudo ground truth masks from readily available images and existing dataset ontologies, by leveraging one more vision foundation model, Segment Anything (SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.]).
At test time, the alignment module projects image patches into text space, and the text-aligned patches are decoded into pixel-wise semantic segmentation features. Each pixel is classified according to its similarity to the text embeddings of a target ontology. This enables the model to use any ontology to produce dataset-specific semantic segmentations on-the-fly. An overview of our method coined as FMbSeg is shown in Fig. 1. Our alignment module is light and fast and with only thousands of unannotated images and few training epochs we achieve SOTA results avoiding the need for expensive pretraining.
@@
The main contributions of this paper are: (1) We propose a method to generate semantic segmentation pseudo annotations with zero pixel or image level labels using CLIP and SAM. (2) We propose composing pretrained language-image and self-supervised vision foundation models by means of a lightweight contrastive alignment module. (3) We apply the composed model to the zeroshot semantic segmentation setting, where we show generalization to never-seen-before semantic segmentation datasets. 4) We achieve state of the art results on annotation-free semantic segmentation task.
@@ 2 Related Work
Recent advancements in large-scale pretraining have led to powerful generalist models, transferable to various tasks and domains  [Bommasani et al.(2021)Bommasani, Hudson, Adeli, Altman, Arora, von Arx, Bernstein, Bohg, Bosselut, Brunskill, et al.]. In particular, expressive vision-language models emerge when scaling contrastive pretraining on text-image pairs, as shown by ALIGN and CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al., Jia et al.(2021)Jia, Yang, Xia, Chen, Parekh, Pham, Le, Sung, Li, and Duerig]. These models have been since applied to a range of vision problems, including open vocabulary semantic segmentation [Ding et al.(2022)Ding, Xue, Xia, and Dai, Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] and image generation [Ramesh et al.(2021)Ramesh, Pavlov, Goh, Gray, Voss, Radford, Chen, and Sutskever]. SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] is a powerful image segmentation model, trained iteratively with weak annotations on one billion images. SAM is not grounded to semantic meaning and although there was an initial attempt by the authors to align SAM with CLIP text embeddings, no quantitative results were reported. DINO [Caron et al.(2021)Caron, Touvron, Misra, Jégou, Mairal, Bojanowski, and Joulin] and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.] are self-supervised image encoders demonstrating high performance in dense prediction tasks. However, DINO models are not aligned with text, making their application to Open Vocabulary scenarios not straightforward. In this work we show an example of foundation models compositionality, where knowledge from different models (CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.] and DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.]) are used to provide strong open vocabulary semantic segmentation model.
@@
Open vocabulary semantic segmentation. Zeroshot semantic segmentation requires a model to provide pixel level labels for unseen categories [Bucher et al.(2019)Bucher, Vu, Cord, and Pérez]. Open vocabulary semantic segmentation generalizes this problem aiming at segmenting an arbitrary set of classes. Zhao et al. (2017) [Zhao et al.(2017)Zhao, Puig, Zhou, Fidler, and Torralba] were the first to introduce the open vocabulary setting by learning a joint embedding for pixels and textual concepts based on WordNet.
@@
CLIP-based methods. The general trend for open vocabulary methods is to build on top of CLIP for better generalization with less supervision [Li et al.(2022b)Li, Zhang, Zhang, Yang, Li, Zhong, Wang, Yuan, Zhang, Hwang, et al., Zhang et al.(2022b)Zhang, Zhang, Hu, Chen, Li, Dai, Wang, Yuan, Hwang, and Gao, Chen et al.(2023)Chen, Li, Lim, Torralba, and Zhao, Ding et al.(2022)Ding, Xue, Xia, and Dai, Xu et al.(2023)Xu, Xiong, Ding, and Tu, Zhong et al.(2022)Zhong, Yang, Zhang, Li, Codella, Li, Zhou, Dai, Yuan, Li, et al., Rao et al.(2022)Rao, Zhao, Chen, Tang, Zhu, Huang, Zhou, and Lu]. LSeg[Li et al.(2022a)Li, Weinberger, Belongie, Koltun, and Ranftl] matches the pixel embeddings to text embedding of CLIP. OpenSeg [Ghiasi et al.(2022)Ghiasi, Gu, Cui, and Lin] learns visual-semantic alignments by aligning each word in a caption to one or few predicted masks. A popular line of work rely on training a class-agnostic mask proposal network, leveraging CLIP embeddings for mask classification with additional techniques to strengthen CLIP image patch embeddings [Xu et al.(2022b)Xu, Zhang, Wei, Lin, Cao, Hu, and Bai, Xu et al.(2023)Xu, Xiong, Ding, and Tu, Han et al.(2023)Han, Zhong, Li, Han, and Ma, Ding et al.(2023)Ding, Wang, and Tu, Yu et al.(2023)Yu, He, Deng, Shen, and Chen]. These works still require a various degrees of pixel-level supervision, which can be expensive to obtain for fine-grained categories. Training on limited available labelled datasets might thus affect their generalization.
@@
Training-free methods. ReCo [Shin et al.(2022)Shin, Xie, and Albanie] builds a dataset of $K$ images for each concept from large-scale unlabelled dataset. Then a nearest neighbour approach is used to produce initial segmentation of a target concept that is then refined by DenseCLIP [Rao et al.(2022)Rao, Zhao, Chen, Tang, Zhu, Huang, Zhou, and Lu]. CLIP-DIY [Wysoczańska et al.(2024)Wysoczańska, Ramamonjisoa, Trzciński, and Siméoni] divides an image into smaller patches that are then classified by CLIP. Patches are then aggregated and transformed to dense prediction using objectness scores from pretrained foreground-background segmentation model. In our work, we rely on SAM to provide accurate masks of detected objects. All training free methods, require a large preprocessing time and usually many steps of refinement, rendering them infeasible. We use the training-free part of our method (section 3.2) as a source of pseudo annotations to produce an efficient model.
@@
Training-based methods without pixel level supervision. Many methods modify the CLIP encoder and update it based on a large set of image-text annotated pairs. GroupVit [Xu et al.(2022a)Xu, De Mello, Liu, Byeon, Breuel, Kautz, and Wang] optimizes a hierarchical pixel grouping strategy integrated in a learned ViT model. TCL [Cha et al.(2023)Cha, Mun, and Roh] trains a decoder to ground masks with language based on datasets of 12M and 3M images. Closer to our work, MaskCLIP [Zhou et al.(2022)Zhou, Loy, and Dai] utilizes CLIP to generate annotations for training a complete segmentation network from scratch. Unlike this work, we train a lightweight module on top of a frozen pretrained vision encoder. SAM-CLIP [Wang et al.(2023b)Wang, Vasu, Faghri, Vemulapalli, Farajtabar, Mehta, Rastegari, Tuzel, and Pouransari] brings semantics to SAM by large-scale fine-tuning and distilation from both SAM and CLIP models with 41M unannotated images. Our work employs SAM as a segmentation teacher and aligns a more expressive vision encoder with CLIP’s pretrained language encoder, hence requiring much less training data and no model retraining. Particularly, our method surpasses previous works performance using only 118 thousand unlabeled images. Our work is the first that aligns off the shelf vision and text encoders, at patches level with little training and a lightweight alignment module; making our methods readily accessible and usable by practitioners for plug and play annotation-free open vocabulary semantic segmentation.
@@ 3 Method
In this work, we deploy 3 foundation models by composition for the task of semantic segmentation: 1) CLIP [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.], pretrained on a million image-text pairs, exhibiting semantic understanding of the image as a whole but not designed for object localization. 2) SAM [Kirillov et al.(2023)Kirillov, Mintun, Ravi, Mao, Rolland, Gustafson, Xiao, Whitehead, Berg, Lo, et al.], trained to segment objects or parts of objects, but lacking proper semantic understanding. 3) DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], producing features that transfer well to many downstream tasks, and that are consistent across similar objects and parts of objects, but without an explicit link to semantic notions. We design a method that leverages the distinctive properties of those foundation models to enable zeroshot and open vocabulary semantic segmentation.
To achieve this, we propose a two-stage approach: first, we leverage CLIP and SAM to generate pseudo semantic masks for a given vocabulary of classes. In the second stage, we use the predictions of the first stage to train a small alignment module that aligns a frozen off-the-shelf image encoder,  DINOv2 [Oquab et al.(2023)Oquab, Darcet, Moutakanni, Vo, Szafraniec, Khalidov, Fernandez, Haziza, Massa, El-Nouby, et al.], with a pretrained text encoder at the patch level, resulting in a strong self-supervised semantic segmentation model.
@@ 3.1 Preliminary
We consider a dataset of images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$ accompanied with a set of categories in the vocabulary $V=\{class_{1},\dots,class_{L}\}$. CLIP’s transformer-based image encoder [Dosovitskiy et al.(2020)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, et al.] takes as input an image ${\mathbf{im}_{i}}\in\mathbb{R}^{C,H,W}$ divided into patches $\mathbf{im}_{i}=[im^{1}_{i},\dots im^{p}_{i},\dots,im^{N}_{i}]$ and extracts a class token $\mathbf{cls}_{i}$ and patches embedding $\mathbf{x}_{i}=[x^{1}_{i},\dots,x^{p}_{i},\dots,x^{N}_{i}]$. CLIP’s text encoder takes as input a text description {a photo of a “$class_{i}$”} and produces the text feature $\mathbf{t}_{i}$, with $class_{i}$ $\in V$ corresponding to the image label. Visual and textual features represented by $\mathbf{cls}_{i}$ and $\mathbf{t}_{i}$ are separately projected to a joint embedding space $\mathbb{R}^{D}$ and the cosine similarity between them is maximized during CLIP’s training. While the original CLIP architecture discards the patch features $\mathbf{x}_{i}$, they can be projected onto the same space $\mathbb{R}^{D}$. This would enable us to compute the similarity of any category in $V$ with individual patches $\mathbf{x}^{p}_{i}$ and produce a rough localization of the objects in the image. Refer to the Appendix for more details on the specific architecture for projecting patch features $\mathbf{x}_{i}$ to $\mathbb{R}^{D}$.

### v3

# Annotation-Free Semantic Segmentation with Vision Foundation Models

@@ Abstract
Semantic Segmentation is one of the most challenging vision tasks, usually requiring large amounts of training data with expensive pixel level annotations. With the success of foundation models and especially vision-language models, recent works attempt to achieve zeroshot semantic segmentation while requiring either large-scale training or additional image/pixel level annotations. In this work, we generate free annotations for any semantic segmentation dataset using existing foundation models. We use CLIP to detect objects and SAM to generate high quality object masks. Next, we build a lightweight module on top of a self-supervised vision encoder, DinoV2, to align the patch features with a pretrained text encoder for zeroshot semantic segmentation. Our approach can bring language-based semantics to any pretrained vision encoder with minimal training, uses foundation models as the sole source of supervision and generalizes from little training data with no annotation.
@@ 1 Introduction
Large-scale and inexpensive training data has recently enabled the surge of foundation models in computer vision [4]. These models have been employed to avoid expensive annotations and computational requirements of many vision tasks [27, 21, 40, 26, 2]. Leveraging foundation models for semantic segmentation requires the model to produce pixel-wise predictions on new datasets, domains and ontologies. Therefore, the model must be 1) promptable with an open set of categories and 2) highly discriminative for dense recognition tasks. Consequently, deploying foundation models to obtain cheap annotation for semantic segmentation is challenging as existing models lack either semantic awareness [40, 24, 21] or local feature robustness [27, 19].
@@
In this paper, we propose a novel approach to open vocabulary semantic segmentation by composition of different foundations models as building blocks and source of supervision. In particular, we employ Contrastive Language-Image Pretraining (CLIP [27]), trained on a large set of image-text pairs, to derive a semantic understanding of different regions within an image [46, 26]. We use Segment Anything (SAM [21]), to supervise the accurate shape and extension of objects detected by CLIP. Next, we train a lightweight module that aligns the generic visual features of a self-supervised task-agnostic foundation model, DinoV2, with the text embedding space of a CLIP model in a contrastive manner. The result is a model that is highly discriminant, generalizable and grounded to semantic meaning and object shape without requiring human-generated segmentation or image annotations.
Figure 1: Overview of our method (FMbSeg). a) Label Generation: We detect object and background categories in an image using a frozen pretrained image-text model (e.g. CLIP, in pink). We select image patches with high similarity to the text representation of the detected categories. We pass the location of those patches to a mask proposal network (e.g. SAM, in green). b) Visual Features Alignment: We use the generated segmentations and detected categories to align features from a more expressive frozen image encoder (e.g. DINOv2, in blue) with a frozen pretrained text encoder. c) Test-Time Inference: At test time, the newly aligned image encoder projects image features into text space. Every pixel is classified according to their similarity to the pre-computed text prototypes of a target ontology.
@@
An overview of our method coined as FMbSeg is shown in Fig. 1. The main contributions of this paper are: (1) We propose a method to generate semantic segmentation pseudo annotations with zero pixel or image level labels using CLIP and SAM. (2) We propose composing pretrained language-image and self-supervised vision foundation models by means of a lightweight contrastive alignment module trained with a uniquely designed loss function. (3) We apply the composed model to the zeroshot semantic segmentation setting, where we show generalization to never-seen-before semantic segmentation datasets. (4) We achieve state of the art results on annotation-free semantic segmentation task.
@@ 2 Related Work
Recent advancements in large-scale pretraining have led to powerful generalist models, transferable to various tasks and domains  [4]. In particular, expressive vision-language models emerge when scaling contrastive pretraining on text-image pairs, as shown by ALIGN and CLIP [27, 19]. These models have been since applied to a range of vision problems, including open vocabulary semantic segmentation [12, 36] and image generation [28]. SAM [21] is a powerful image segmentation model, trained iteratively with weak annotations on one billion images. SAM is not grounded to semantic meaning and although there was an initial attempt by the authors to align SAM with CLIP text embeddings, no quantitative results were reported. DINO [7] and DINOv2 [26] are self-supervised image encoders demonstrating high performance in dense prediction tasks. However, DINO models are not aligned with text, making their application to Open Vocabulary scenarios not straightforward. In this work we show an example of foundation models compositionality, where knowledge from different models (CLIP [27], SAM [21] and DINOv2 [26]) are incorporated to build an open vocabulary semantic segmentation model.
@@
Open vocabulary semantic segmentation. Zeroshot semantic segmentation requires a model to provide pixel level labels for unseen categories [5]. Open vocabulary semantic segmentation generalizes this problem aiming at segmenting an arbitrary set of classes. Zhao et al. [42] first introduced the open vocabulary setting by learning a joint embedding for pixels and textual concepts based on WordNet.
@@
CLIP-based methods. The general trend for open vocabulary methods is to build on top of CLIP for better generalization with less supervision [23, 41, 10, 12, 38, 44, 29]. LSeg[22] matches the pixel embeddings to text embedding of CLIP. OpenSeg [16] learns visual-semantic alignments by aligning each word in a caption to one or few predicted masks. A popular line of work relies on training a class-agnostic mask proposal network, leveraging CLIP embeddings for mask classification with additional techniques to strengthen CLIP image patch embeddings [37, 38, 17, 13, 39]. These works still require various degrees of pixel-level supervision, which can be expensive to obtain for fine-grained categories.
@@
Training-free methods. ReCo [30] builds a dataset of $K$ images for each concept from a large-scale unlabelled dataset. Then a nearest neighbour approach is used to produce initial segmentation of a target concept that is then refined by DenseCLIP [29]. CLIP-DIY [34] divides an image into smaller patches that are then classified by CLIP. Patches are then aggregated and transformed to dense prediction using objectness scores from pretrained foreground-background segmentation model. In our work, we rely on SAM to provide accurate masks of detected objects. All training free methods, require a large preprocessing time and usually many steps of refinement, rendering them infeasible. We use the training-free part of our method (section 3.2) as a source of pseudo annotations to produce an efficient model.
@@
Training-based methods without pixel level supervision. Many methods modify the CLIP encoder and update it based on a large set of image-text annotated pairs. GroupVit [36] optimizes a hierarchical pixel grouping strategy integrated in a learned ViT model. TCL [8] trains a decoder to ground masks with language based on datasets of 12M and 3M images. ZeroSeg [9] distills localized semantic information from multi-scale views to a segmentation model via different loss functions. Closer to our work, MaskCLIP [46] utilizes CLIP to generate annotations for training a complete segmentation network from scratch. These works are restricted to CLIP due to its language capabilities and hence require large scale training to create discriminate patch-level features. To the contrary, we map DINOv2 [26] accurate patch features to CLIP text space with minimal training. SAM-CLIP [33] brings semantics to SAM by large-scale fine-tuning and distillation from both SAM and CLIP models with 41M unannotated images. Our work only employs SAM as a segmentation teacher and aligns a more expressive vision encoder with CLIP’s pretrained language encoder, hence requiring much less training data and no model retraining. Particularly, our method surpasses previous works performance using only 118 thousand unlabeled images. Our work is the first to align off-the-shelf vision and text encoders at the patch level with minimal training and a lightweight alignment module, making our method readily accessible for annotation-free semantic segmentation and usable by practitioners for plug and play semantic segmentation methods on various domains.
@@ 3 Method
In this work, we deploy 3 foundation models by composition for the task of semantic segmentation: 1) CLIP [27], pretrained on a million image-text pairs, exhibiting semantic understanding of the image as a whole but not designed for object localization. 2) SAM [21], trained to segment objects or parts of objects, but lacking proper semantic understanding. 3) DINOv2 [26], producing features that transfer well to many downstream tasks, and that are consistent across similar objects and parts of objects, but without an explicit link to semantic notions. We design a method that leverages the distinctive properties of those foundation models to enable zeroshot and open vocabulary semantic segmentation.
To achieve this, we propose a two-stage approach: first, we leverage CLIP and SAM to generate pseudo semantic masks for a given vocabulary of classes. In the second stage, we use the predictions of the first stage to train a small alignment module that aligns a frozen off-the-shelf image encoder,  DINOv2 [26], with a pretrained text encoder at the patch level, resulting in a strong self-supervised semantic segmentation model.
@@ 3.1 Preliminary
We consider a dataset of images $\mathbf{D}=\{im_{i}\}^{M}_{i=1}$ accompanied with a set of categories in the vocabulary $V=\{class_{1},\dots,class_{k},\dots,class_{K}\}$. CLIP’s transformer-based image encoder [14] takes as input an image ${\mathbf{im}_{i}}\in\mathbb{R}^{C,H,W}$ divided into patches $\mathbf{im}_{i}=[im^{1}_{i},\dots im^{p}_{i},\dots,im^{N}_{i}]$ and extracts a class token $\mathbf{cls}_{i}$ and patches embedding $\mathbf{x}_{i}=[x^{1}_{i},\dots,x^{p}_{i},\dots,x^{N}_{i}]$. CLIP’s text encoder takes as input a text description {a photo of a “$class_{k}$”} and produces the text feature $\mathbf{t}_{k}$, with $class_{k}$ $\in V$ corresponding to the image label. Visual and textual features represented by $\mathbf{cls}_{i}$ and $\mathbf{t}_{k}$ are separately projected to a joint embedding space $\mathbb{R}^{D}$ and the cosine similarity between them is maximized during CLIP’s training. While the original CLIP architecture discards the patch features $\mathbf{x}^{p}_{i}$, they can be projected onto the same space $\mathbb{R}^{D}$. This would enable us to compute the similarity of any category in $V$ with individual patches $\mathbf{x}^{p}_{i}$ and produce a rough localization of the objects in the image. Refer to the Appendix for more details on the specific architecture for projecting patch features $\mathbf{x}^{p}_{i}$ to $\mathbb{R}^{D}$.